Calculate the percentage purity of CaCO3 when 50g of marble reacts with an excess of HCl to make 15 g of CO2.


1
Expert's answer
2021-01-04T03:47:36-0500

"CaCO_3+2HCl\u2192CaCl_2+H_2O+CO_2\n\u200b"


From the reaction above, 1 mole of CO2 is produced by 1 mole of CaCO3.

This means that, 44g of CO2 is produced by 100g of CaCO3.


15g of CO2 is therefore produced by "\frac{15×100}{44}"g of CaCO3.

34g of CaCO3 produces 15g of CO2.


Percentage of purity = "\dfrac{\textsf{Mass of pure sample}}{\textsf{Total mass of impure sample}} ×100\%= \dfrac{34}{50}×100\%=68\%"


Therefore, the percentage purity of the marble is 68%.